# RAPTOR: Ridge-Adaptive Logistic Probes

### v1

@@ Abstract
Large Language Models, Interpretability, Concept Steering, Logistic Regression, High-dimensional Asymptotics
@@ 3.1 Ridge-Adaptive Logistic Probe
$\mu_{j}^{(\ell)}$ $=\frac{1}{|\mathcal{I}_{\mathrm{tr}}|}\sum_{i\in\mathcal{I}_{\mathrm{tr}}}h_{ij}^{(\ell)},$ $j$ $=1,\dots,p,$ (2)
$s_{j}^{(\ell)}$ $=\sqrt{\frac{1}{|\mathcal{I}_{\mathrm{tr}}|}\sum_{i\in\mathcal{I}_{\mathrm{tr}}}\big(h_{ij}^{(\ell)}-\mu_{j}^{(\ell)}\big)^{2}},$ $j$ $=1,\dots,p,$ (3)
where $s^{(\ell)}\in\mathbb{R}^{p}$ is the elementwise standard deviation (zero entries are replaced by $1$ in implementation). We form standardized features $x_{ij}^{(\ell)}=\frac{h_{ij}^{(\ell)}-\mu_{j}^{(\ell)}}{s_{j}^{(\ell)}},j=1,\dots,p,$ and apply the same transform to validation and test embeddings. This train-only standardization prevents information leakage and stabilizes optimization by controlling feature scales, so that ridge tuning is not dominated by a small subset of large-variance coordinates.
@@
$\mathcal{L}_{\lambda}^{(\ell)}(w,b;\mathcal{I})\;=$ $\;\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\log\!\Big(1+\exp\!\big(-\tilde{y}_{i}(w^{\top}x_{i}^{(\ell)}+b)\big)\Big)$ (4)
$\;+\;\frac{\lambda}{2}\|w\|_{2}^{2}$
@@
After fitting in standardized coordinates, we fold parameters back to the original embedding space so that the resulting concept vector can be injected into native (unstandardized) layer representations: $\hat{\omega}^{(\ell)}_{j}=\frac{\hat{w}^{(\ell)}_{j}}{s_{j}^{(\ell)}},j=1,\dots,p,$ $\hat{b}_{\mathrm{orig}}^{(\ell)}=\hat{b}^{(\ell)}-\langle\hat{\omega}^{(\ell)},\mu^{(\ell)}\rangle.$
@@ 3.2 Steering Setup
Given a learned layer-wise concept direction, additive steering only requires choosing the injection strength. We adopt the GCAV per-sample calibration rule (Zhang et al., 2025; Xu et al., 2024) as an off-the-shelf way to set this strength without retraining.
@@ 5.3 Fixed-point characterization
$\eta_{\gamma}(u)$ $=\operatorname{prox}_{\gamma\rho}(u)=\operatorname*{arg\,min}_{t\in\mathbb{R}}\Big\{\rho(t)+\frac{1}{2\gamma}(t-u)^{2}\Big\},$ (14)
$\qquad\gamma>0.$
@@ 5.4 Deterministic limit of test accuracy
$\begin{split}\mathrm{Acc}(\delta,\lambda,\kappa)=\mathbb{E}_{Z}\Big[&\ \sigma(\kappa Z)\,\Phi\!\left(\frac{\kappa\bar{\alpha}}{\bar{\sigma}}Z\right)\\ +&\ \sigma(-\kappa Z)\left(1-\Phi\!\left(\frac{\kappa\bar{\alpha}}{\bar{\sigma}}Z\right)\right)\Big]\end{split}$ (20)
@@ 6 Discussion and Conclusion
This appendix provides a self-contained CGMT derivation of Theorem 5.1. After matching notation, the resulting fixed-point equations are equivalent to those in Salehi et al. (2019). The main text states the final fixed-point system (Theorem 5.1) and the limiting test accuracy (Proposition 5.2); here we give a CGMT-based route from the ERM (12) to the deterministic characterization.
@@ A.1.2 Estimator (ridge logistic regression)
We consider the ridge-logistic scaling used in Salehi et al. (2019):
@@ A.2 From ERM to a convex min–max (Primary Optimization)
$\boxed{\begin{aligned} \Phi(\mathbf{M},\mathbf{y}):=\min_{\mathbf{z}\in\mathbb{R}^{p}}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{&-\mathbf{w}^{\top}\operatorname{diag}(\mathbf{y})\mathbf{M}\mathbf{z}\\ &-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}\|\mathbf{z}\|^{2}\Big\}\end{aligned}}$ (26)
@@ A.3.2 Decompose both the matrix and the primal variable
$\Phi$ $=\min_{\alpha\in\mathbb{R},\ \mathbf{z}_{\perp}\in\mathbf{v}^{\perp}}\max_{\mathbf{w}\in[0,1]^{n}}\Big\{\$(29)
$\quad-\alpha\,\mathbf{w}^{\top}\operatorname{diag}(\mathbf{y})\mathbf{u}$
$-\mathbf{w}^{\top}\operatorname{diag}(\mathbf{y})\mathbf{M}_{\perp}\mathbf{z}_{\perp}$
$-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}\big(\alpha^{2}+\|\mathbf{z}_{\perp}\|^{2}\big)\Big\}.$
@@
$\Phi$ $=\min_{\alpha\in\mathbb{R},\ \mathbf{z}_{\perp}\in\mathbf{v}^{\perp}}\max_{\mathbf{w}\in[0,1]^{n}}\Big\{\$(30)
$\quad-\alpha\,\mathbf{w}^{\top}\mathbf{s}$
$-\mathbf{w}^{\top}\mathbf{A}\mathbf{z}_{\perp}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})$
$+\frac{n\lambda}{2}\big(\alpha^{2}+\|\mathbf{z}_{\perp}\|^{2}\big)\Big\}.$
@@ A.4.1 CGMT form used
We invoke the convex Gaussian min–max theorem (CGMT) in the standard bilinear form; see, e.g., Thrampoulidis et al. (2015, 2018).
@@
$\qquad\Phi_{\mathrm{AO}}$ $:=\min_{\mathbf{z}_{\perp}\in\mathcal{S}}\max_{\mathbf{w}\in\mathcal{T}}\ \|\mathbf{z}_{\perp}\|\,\mathbf{g}^{\top}\mathbf{w}+\|\mathbf{w}\|\,\mathbf{h}^{\top}\mathbf{z}_{\perp}+$
$\quad\Psi(\mathbf{z}_{\perp},\mathbf{w}).$
@@ A.4.2 Apply CGMT to (30)
$\phi$ $=\min_{\alpha\in\mathbb{R},\ \mathbf{z}_{\perp}\in\mathbf{v}^{\perp}}\max_{\mathbf{w}\in[0,1]^{n}}\Big\{\$(31)
$\quad-\alpha\,\mathbf{w}^{\top}\mathbf{s}+\|\mathbf{z}_{\perp}\|\,\mathbf{g}^{\top}\mathbf{w}$
$\quad+\|\mathbf{w}\|\,\mathbf{h}^{\top}\mathbf{z}_{\perp}$
$\quad-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}\big(\alpha^{2}+\|\mathbf{z}_{\perp}\|^{2}\big)\Big\}.$
@@ A.5.1 Eliminate the direction of $\mathbf{z}_{\perp}$
$\phi=$ $\min_{\alpha\in\mathbb{R},\ r\geq 0}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{-\alpha\,\mathbf{w}^{\top}\mathbf{s}$ (32)
$\quad+r\,\mathbf{g}^{\top}\mathbf{w}-r\,\|\mathbf{h}\|\,\|\mathbf{w}\|$
$\quad-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}(\alpha^{2}+r^{2})\Big\}.$
@@ A.5.2 Two variational identities used to decouple norms and squares
$\phi$ $=\min_{\alpha,\ r\geq 0}\ \max_{\tau>0}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{-\alpha\,\mathbf{w}^{\top}\mathbf{s}$ (33)
$\quad+r\,\mathbf{g}^{\top}\mathbf{w}-\frac{\tau}{2}\|\mathbf{w}\|^{2}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})$
$\quad+\frac{n\lambda}{2}(\alpha^{2}+r^{2})-\frac{r^{2}\|\mathbf{h}\|^{2}}{2\tau}\Big\}.$
@@ A.5.4 Eliminate $\alpha$ explicitly
$\phi$ $=\min_{r\geq 0}\ \max_{\tau>0}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{r\,\mathbf{g}^{\top}\mathbf{w}$ (34)
$\quad-\frac{\tau}{2}\|\mathbf{w}\|^{2}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}r^{2}$
$\quad-\frac{r^{2}\|\mathbf{h}\|^{2}}{2\tau}-\frac{(\mathbf{w}^{\top}\mathbf{s})^{2}}{2n\lambda}\Big\}.$
@@ A.5.5 Decouple $(\mathbf{w}^{\top}\mathbf{s})^{2}$
$\phi$ $=\min_{r\geq 0}\ \max_{\tau>0}\ \min_{\gamma\in\mathbb{R}}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{\sum_{i=1}^{n}\Big[(rg_{i}-\gamma s_{i})w_{i}$ (35)
$\quad-\frac{\tau}{2}w_{i}^{2}-\rho^{\!*}(w_{i})\Big]+\frac{n\lambda}{2}r^{2}-\frac{r^{2}\|\mathbf{h}\|^{2}}{2\tau}$
$\quad+\frac{n\lambda}{2}\gamma^{2}\Big\}.$
@@ A.7 Deterministic fixed-point characterization
At this point, the AO reduces (after substituting the coordinate maximizers) to a scalar saddle problem whose empirical averages concentrate. The resulting limiting KKT conditions can be expressed as a closed system for the geometric parameters of $\hat{\mathbf{z}}$. We record the final characterization in the notation used below; see Salehi et al. (2019) for derivation details under standard CGMT regularity conditions.
@@ A.8.2 Closed form for accuracy of the zero-threshold classifier
$\mathrm{Acc}$ $=\mathbb{E}_{Z}\Big[\ \sigma(\kappa Z)\,\Phi\!\Big(\frac{\bar{\alpha}\kappa Z}{\bar{\sigma}}\Big)$ (40)
$+\sigma(-\kappa Z)\,\Big(1-\Phi\!\Big(\frac{\bar{\alpha}\kappa Z}{\bar{\sigma}}\Big)\Big)\Big]$
@@ B.3 Steering protocol and hyperparameters
Rather than choosing a fixed injection strength, we determine the minimal per-layer strength that achieves an extreme probe target (Zhang et al., 2025; Xu et al., 2024). Fix a layer $\ell$ and let $h^{(\ell)}\in\mathbb{R}^{p}$ denote the pre-intervention hidden state. We steer by an additive intervention (Eq. 1 in the main text),
@@
$\alpha_{\ell}\;=\;\arg\min_{\alpha\geq 0}$ $|\alpha|$ (43)
s.t. $P_{m}\!\left(h^{(\ell)}+\alpha v_{\ell}\right)\;\geq\;p_{\star},$

### v2

@@ Abstract
Keywords large language models $\cdot$ activation steering $\cdot$ representation steering $\cdot$ concept vectors $\cdot$ probing, interpretability $\cdot$ robustness $\cdot$ logistic probes $\cdot$ ridge regularization
@@ 3.1 Ridge-Adaptive Logistic Probe
$\mu_{j}^{(\ell)}=\frac{1}{|\mathcal{I}_{\mathrm{tr}}|}\sum_{i\in\mathcal{I}_{\mathrm{tr}}}h_{ij}^{(\ell)},\qquad j=1,\dots,p,$ (2)
$s_{j}^{(\ell)}=\sqrt{\frac{1}{|\mathcal{I}_{\mathrm{tr}}|}\sum_{i\in\mathcal{I}_{\mathrm{tr}}}\big(h_{ij}^{(\ell)}-\mu_{j}^{(\ell)}\big)^{2}},\qquad j=1,\dots,p,$ (3)
where $s^{(\ell)}\in\mathbb{R}^{p}$ is the elementwise standard deviation (zero entries are replaced by $1$ in implementation). We form standardized features
$x_{ij}^{(\ell)}=\frac{h_{ij}^{(\ell)}-\mu_{j}^{(\ell)}}{s_{j}^{(\ell)}},\qquad j=1,\dots,p,$ (4)
and apply the same transform to validation and test embeddings. This train-only standardization prevents information leakage and stabilizes optimization by controlling feature scales, so that ridge tuning is not dominated by a small subset of large-variance coordinates.
@@
$\mathcal{L}_{\lambda}^{(\ell)}(w,b;\mathcal{I})\;=$ $\;\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\log\!\Big(1+\exp\!\big(-\tilde{y}_{i}(w^{\top}x_{i}^{(\ell)}+b)\big)\Big)\;+\;\frac{\lambda}{2}\|w\|_{2}^{2}$ (5)
@@
After fitting in standardized coordinates, we fold parameters back to the original embedding space so that the resulting concept vector can be injected into native (unstandardized) layer representations:
$\hat{\omega}^{(\ell)}_{j}=\frac{\hat{w}^{(\ell)}_{j}}{s_{j}^{(\ell)}},j=1,\dots,p,\qquad\hat{b}_{\mathrm{orig}}^{(\ell)}=\hat{b}^{(\ell)}-\langle\hat{\omega}^{(\ell)},\mu^{(\ell)}\rangle$
@@ 3.2 Steering Setup
Given a learned layer-wise concept direction, additive steering only requires choosing the injection strength. We adopt the GCAV per-sample calibration rule Zhang et al. (2025); Xu et al. (2024) as an off-the-shelf way to set this strength without retraining.
@@ 5.3 Fixed-point characterization
$\eta_{\gamma}(u)$ $=\operatorname{prox}_{\gamma\rho}(u)=\operatorname*{arg\,min}_{t\in\mathbb{R}}\Big\{\rho(t)+\frac{1}{2\gamma}(t-u)^{2}\Big\},\qquad\gamma>0.$ (15)
@@ 5.4 Deterministic limit of test accuracy
$\begin{split}\mathrm{Acc}(\delta,\lambda,\kappa)=\mathbb{E}_{Z}\Big[\ \sigma(\kappa Z)\,\Phi\!\left(\frac{\kappa\bar{\alpha}}{\bar{\sigma}}Z\right)+\ \sigma(-\kappa Z)\left(1-\Phi\!\left(\frac{\kappa\bar{\alpha}}{\bar{\sigma}}Z\right)\right)\Big]\end{split}$ (21)
@@ 6 Discussion and Conclusion
This appendix provides a self-contained CGMT derivation of Theorem 5.1. After matching notation, the resulting fixed-point equations are equivalent to those in Salehi et al. [2019]. The main text states the final fixed-point system (Theorem 5.1) and the limiting test accuracy (Proposition 5.2); here we give a CGMT-based route from the ERM (13) to the deterministic characterization.
@@ A.1.2 Estimator (ridge logistic regression)
We consider the ridge-logistic scaling used in Salehi et al. [2019]:
@@ A.2 From ERM to a convex min–max (Primary Optimization)
$\boxed{\Phi(\mathbf{M},\mathbf{y}):=\min_{\mathbf{z}\in\mathbb{R}^{p}}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{-\mathbf{w}^{\top}\operatorname{diag}(\mathbf{y})\mathbf{M}\mathbf{z}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}\|\mathbf{z}\|^{2}\Big\}}$ (27)
@@ A.3.2 Decompose both the matrix and the primal variable
$\Phi=\min_{\alpha\in\mathbb{R},\ \mathbf{z}_{\perp}\in\mathbf{v}^{\perp}}\max_{\mathbf{w}\in[0,1]^{n}}\Big\{\ \quad\alpha\,\mathbf{w}^{\top}\operatorname{diag}(\mathbf{y})\mathbf{u}-\mathbf{w}^{\top}\operatorname{diag}(\mathbf{y})\mathbf{M}_{\perp}\mathbf{z}_{\perp}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}\big(\alpha^{2}+\|\mathbf{z}_{\perp}\|^{2}\big)\Big\}.$ (30)
@@
$\Phi=\min_{\alpha\in\mathbb{R},\ \mathbf{z}_{\perp}\in\mathbf{v}^{\perp}}\max_{\mathbf{w}\in[0,1]^{n}}\Big\{\ \\ \quad-\alpha\,\mathbf{w}^{\top}\mathbf{s}\\ -\mathbf{w}^{\top}\mathbf{A}\mathbf{z}_{\perp}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})\\ +\frac{n\lambda}{2}\big(\alpha^{2}+\|\mathbf{z}_{\perp}\|^{2}\big)\Big\}.$ (31)
@@ A.4.1 CGMT form used
We invoke the convex Gaussian min–max theorem (CGMT) in the standard bilinear form; see, e.g., Thrampoulidis et al. [2015, 2018].
@@
$\qquad\Phi_{\mathrm{AO}}$ $:=\min_{\mathbf{z}_{\perp}\in\mathcal{S}}\max_{\mathbf{w}\in\mathcal{T}}\ \|\mathbf{z}_{\perp}\|\,\mathbf{g}^{\top}\mathbf{w}+\|\mathbf{w}\|\,\mathbf{h}^{\top}\mathbf{z}_{\perp}+\quad\Psi(\mathbf{z}_{\perp},\mathbf{w}).$
@@ A.4.2 Apply CGMT to (31)
$\phi$ $=\min_{\alpha\in\mathbb{R},\ \mathbf{z}_{\perp}\in\mathbf{v}^{\perp}}\max_{\mathbf{w}\in[0,1]^{n}}\Big\{\ -\alpha\,\mathbf{w}^{\top}\mathbf{s}+\|\mathbf{z}_{\perp}\|\,\mathbf{g}^{\top}\mathbf{w}+\|\mathbf{w}\|\,\mathbf{h}^{\top}\mathbf{z}_{\perp}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}\big(\alpha^{2}+\|\mathbf{z}_{\perp}\|^{2}\big)\Big\}.$ (32)
@@ A.5.1 Eliminate the direction of $\mathbf{z}_{\perp}$
$\phi=\min_{\alpha\in\mathbb{R},\ r\geq 0}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{-\alpha\,\mathbf{w}^{\top}\mathbf{s}+r\,\mathbf{g}^{\top}\mathbf{w}-r\,\|\mathbf{h}\|\,\|\mathbf{w}\|-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}(\alpha^{2}+r^{2})\Big\}.$ (33)
@@ A.5.2 Two variational identities used to decouple norms and squares
$\phi$ $=\min_{\alpha,\ r\geq 0}\ \max_{\tau>0}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{-\alpha\,\mathbf{w}^{\top}\mathbf{s}+r\,\mathbf{g}^{\top}\mathbf{w}-\frac{\tau}{2}\|\mathbf{w}\|^{2}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}(\alpha^{2}+r^{2})-\frac{r^{2}\|\mathbf{h}\|^{2}}{2\tau}\Big\}.$ (34)
@@ A.5.4 Eliminate $\alpha$ explicitly
$\phi$ $=\min_{r\geq 0}\ \max_{\tau>0}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{r\,\mathbf{g}^{\top}\mathbf{w}-\frac{\tau}{2}\|\mathbf{w}\|^{2}-\sum_{i=1}^{n}\rho^{\!*}(w_{i})+\frac{n\lambda}{2}r^{2}-\frac{r^{2}\|\mathbf{h}\|^{2}}{2\tau}-\frac{(\mathbf{w}^{\top}\mathbf{s})^{2}}{2n\lambda}\Big\}.$ (35)
@@ A.5.5 Decouple $(\mathbf{w}^{\top}\mathbf{s})^{2}$
$\phi=\min_{r\geq 0}\ \max_{\tau>0}\ \min_{\gamma\in\mathbb{R}}\ \max_{\mathbf{w}\in[0,1]^{n}}\Big\{\sum_{i=1}^{n}\Big[(rg_{i}-\gamma s_{i})w_{i}-\frac{\tau}{2}w_{i}^{2}-\rho^{\!*}(w_{i})\Big]+\frac{n\lambda}{2}r^{2}-\frac{r^{2}\|\mathbf{h}\|^{2}}{2\tau}+\frac{n\lambda}{2}\gamma^{2}\Big\}.$ (36)
@@ A.7 Deterministic fixed-point characterization
At this point, the AO reduces (after substituting the coordinate maximizers) to a scalar saddle problem whose empirical averages concentrate. The resulting limiting KKT conditions can be expressed as a closed system for the geometric parameters of $\hat{\mathbf{z}}$. We record the final characterization in the notation used below; see Salehi et al. [2019] for derivation details under standard CGMT regularity conditions.
@@ A.8.2 Closed form for accuracy of the zero-threshold classifier
$\mathrm{Acc}=\mathbb{E}_{Z}\Big[\ \sigma(\kappa Z)\,\Phi\!\Big(\frac{\bar{\alpha}\kappa Z}{\bar{\sigma}}\Big)+\sigma(-\kappa Z)\,\Big(1-\Phi\!\Big(\frac{\bar{\alpha}\kappa Z}{\bar{\sigma}}\Big)\Big)\Big]$ (41)
@@ B.3 Steering protocol and hyperparameters
Rather than choosing a fixed injection strength, we determine the minimal per-layer strength that achieves an extreme probe target Zhang et al. [2025], Xu et al. [2024]. Fix a layer $\ell$ and let $h^{(\ell)}\in\mathbb{R}^{p}$ denote the pre-intervention hidden state. We steer by an additive intervention (Eq. 1 in the main text),
@@
$\alpha_{\ell}\;=\;\arg\min_{\alpha\geq 0}\;|\alpha|\quad\text{s.t.}P_{m}\!\left(h^{(\ell)}+\alpha v_{\ell}\right)\;\geq\;p_{\star},$ (44)